A sealed balloon is filled with 1.25L of helium at 23.0c and 1.00bar. The balloon rises to a point in the atmosphere the pressure is 0.250 bar and the temperature is-31.0c. What is the volume of the balloon under these new conditions?
1
Expert's answer
2020-03-23T10:15:11-0400

"V_1=1.25L"

"P_1=1 bar"

"T_1=23\u00b0C=296K"

"P_2\u200b=0.250bar"

"T_2=-31\u00b0C=244K"

"\frac{P_2V_2}{T_2}=\frac{P_1V_1}{T_1}"

"V_2=\frac{P_1V_1T_2}{T_1P_2}"

"V_2=\frac{1×1.25×244}{296×0.25}"

"V_2=4.122L"